Perform all indicated operations and write the answer with positive integer exponents


(b^-1 + c^ -1) / (bc)-1


1
Expert's answer
2021-03-07T17:24:12-0500

"((b^{-1}+c^{-1}))/(bc)^{-1}=\frac{\frac{1}{b}+\frac{1}{c}}{\frac{1}{bc}} =\frac{\frac{c}{bc}+\frac{b}{bc}}{\frac{1}{bc}}=\frac{\frac{(c+b)}{bc}}{\frac{1}{bc}}=\frac{c+b}{bc}∙\frac{bc}{1}=\frac{c+b}{1}=c+b"